I am working on the Zero Product Property
This problem is confusing. How do factor where I left off?
3x^2 + 16x = -5
3x^2 + 16x +5 = 0

Answer by mananth(16949)

3x^2 + 16x = -5
3x^2 + 16x +5 = 0
3x^2+15x+x+5=0
3x(x+5)+1(x+5)=0
(x+5)(3x+1)=0